OpenStudy (anonymous):

Please help me solve this..if the point (8,-4) is on the graph of y=f(x). find the corresponding point on the graph of g(x)=f(4)

jimthompson5910 (jim_thompson5910):

How are f(x) and g(x) connected?

OpenStudy (anonymous):

Find the corresponding point on the graph of g(x)=f(4x)

jimthompson5910 (jim_thompson5910):

Does it provide you a graph?

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

I might be wrong, but it seems like something is missing. I'm not sure what though.

jimthompson5910 (jim_thompson5910):

hmm g(x) = f(4x)

OpenStudy (anonymous):

I'm kind of confuse myself, don't even know where to start

jimthompson5910 (jim_thompson5910):

(8,-4) lies on f(x)

jimthompson5910 (jim_thompson5910):

so x = 8, y = -4

jimthompson5910 (jim_thompson5910):

g(x) = f(4x) g(x) = f(4*8) g(x) = f(32) then I get stuck. So I'm probably going the wrong way

jimthompson5910 (jim_thompson5910):

or g(x) = f(4x) g(2) = f(4*2) ... plug in x = 2 g(2) = f(8) g(2) = -4

jimthompson5910 (jim_thompson5910):

ok yeah I figured it out

jimthompson5910 (jim_thompson5910):

if you were to plug in x = 2 into g(x) = f(4x), you get g(2) = -4 So that means when x = 2, g(x) = -4 therefore, we know for sure the point (2,-4) lies on g(x)

OpenStudy (anonymous):

And why using x=2

jimthompson5910 (jim_thompson5910):

well we know for sure that f(8) = -4 this is given from the point (8,-4) that lies on f(x)

jimthompson5910 (jim_thompson5910):

then we're given g(x) = f(4x) so I want to make that 4x into 8 that happens when x = 2

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

you kinda have to start with what you have, and with what you want and try to find a midway meeting point or start with what you want and work backwards to connect it to what you have

jimthompson5910 (jim_thompson5910):

that's one way to think of it

OpenStudy (anonymous):

Makes sense, thanks.

jimthompson5910 (jim_thompson5910):

you're welcome